Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 100 miles per hour and train B is traveling at 104 miles per hour. Train A passes a station at 9:25 PM. If B passes the same station at 9:55 P.M at what time will train B catch up to A?
Answer by checkley77(12844)

D=RT
D=100T FOR TRAIN A.
D=104(T-.5) FOR TRAIN B.
THE DISTANCES ARE THE SAME SO SET THE 2 EQUATIONS EQUAL & SOLVE FOR T.
100T=104(T-.5)
100T=104T-52
100T-104T=-52
-4T=-52
T=-52/-4
T=13 HOURS AFTER TRAIN A LEAVES THE STATION TRAIN B WILL CATCH TRAIN A.
PROOF:
100*13=104(13-.5)
1,300=1,352-52
1,300=1,300